# On Adversarial Attacks In Acoustic Drone Localization

### v1

@@ Abstract
Multi-rotor aerial autonomous vehicles (MAVs, more widely known as ”drones”) have been generating increased interest in recent years due to their growing applicability in a vast and diverse range of fields (e.g., agriculture, commercial delivery, search and rescue). The sensitivity of visual-based methods to lighting conditions and occlusions had prompted growing study of navigation reliant on other modalities, such as acoustic sensing. A major concern in using drones in scale for tasks in non-controlled environments is the potential threat of adversarial attacks over their navigational systems, exposing users to mission-critical failures, security breaches, and compromised safety outcomes that can endanger operators and bystanders. While previous work shows impressive progress in acoustic-based drone localization, prior research in adversarial attacks over drone navigation only addresses visual sensing-based systems. In this work, we aim to compensate for this gap by supplying a comprehensive analysis of the effect of PGD adversarial attacks over acoustic drone localization. We furthermore develop an algorithm for adversarial perturbation recovery, capable of markedly diminishing the affect of such attacks in our setting. The code for reproducing all experiments will be released upon publication.
@@ 1 Introduction
Recent years have brought tremendous progress in the use of multi-rotor autonomous aerial vehicles (drones) across a wide range of industries, with their applications expanding rapidly in both scope and variety (Ayamga et al., 2021; Merkert & Bushell, 2020; Moshref-Javadi & Winkenbach, 2021). These industries encompass humanity-crucial applications, such as emergency medicine (Johnson et al., 2021; Zailani et al., 2020), sustainability (Dutta & Goswami, 2020; Mahroof et al., 2021) and disaster control (Daud et al., 2022; Ishiwatari, 2024), as well as life-quality and recreational uses (Tan et al., 2021; Benarbia & Kyamakya, 2021). One major contributor to the growing popularity of drones would be their potential in reducing the risks and added costs often incurred with human involvement upon operating in challenging, unfamiliar and unpredictable environments.
@@
The successful deployment of drones, or any non-stationary agent operating in an environment, heavily relies on the ability to efficiently and accurately navigate within it. This observation has fostered remarkable progress in drone navigation and localization in recent years (Kang et al., 2015; Dijkshoorn, 2012; Sorbelli et al., 2018), expedited by the tremendous progress achieved with the advent of deep-learning (Yousaf et al., 2022; Bisio et al., 2021; Zhang et al., 2022). The currently prevalent approaches for drone navigation mostly rely on either GPS combined with inertial systems, passive visual odometry, active range sensors like LiDAR, or some combination thereof (Arafat et al., 2023; Aburaya et al., 2024; Niu et al., 2022; Debeunne & Vivet, 2020). Using such modalities would be a natural choice in many cases, given their accuracy, availability, and vast amounts of relevant existing research. Nonetheless, in many common scenarios, the reliance on such methods is either challenging (Arafat et al., 2023; Al-Radaideh & Sun, 2021), ineffective (Dreissig et al., 2023), or altogether infeasible (Al-Radaideh & Sun, 2021; Meles et al., 2023). Such difficulties have encouraged growing research in localization relying on other modalities, with one of the most prominent approaches being acoustic localization (He et al., 2023; Sun et al., 2023; Shor et al., 2024). These methods rely on acoustic signals (solely or jointly with other modalities) to perform localization more robust (Famili et al., 2022) to the various obstructions hindering the use of previously mentioned counterparts.
Alongside the growing prevalence of drones in our lives, there also comes growing concern of increased exposure to potential adversarial attacks. As previous studies indicate (Mynuddin et al., 2023; Wisniewski et al., 2024; Guesmi & Shafique, 2024), adversarial attacks can be used to manipulate drones into undesirable or harmful behavior, and even endanger bystanders (e.g. adversarial crashes (Fu et al., 2021)). While past work does tackle defense against such attacks (Mynuddin et al., 2024; Wang et al., 2023), these studies only focus on adversarial attacks over visual or LiDAR channels. Despite of the rapidly increasing popularity acoustic drone localization is garnering, to the best of our knowledge, no prior work has investigated adversarial attacks and defenses related to acoustic sensing-led navigation.
To bridge this gap, in this work we provide, to the best of our knowledge for the first time, the formulation and impact analysis of white-box adversarial attacks over acoustic drone localization. In our setting we study adversarial sound generated by a single stationary source, external to the drone. Furthermore, we rely on the phase modulation mechanism developed in the propulsion noise-based localization algorithm proposed in (Shor et al., 2024) to initiate a novel adversarial defense algorithm, capable of not only increasing adversarial robustness, but also delineate the sampled perturbation from the ”clean” input signal, assuming only minimal assumptions over the attack source. We also expand the algorithm from (Shor et al., 2024) to perform localization over acoustic signals recorded from actual complex acoustic settings, rather than from fully-simulated data, as presented in the original work. We emphasize that while our work focuses on drone localization, it is applicable to any acoustic localization tasks involving a single autonomous agent.
In summary, our contribution is fourfold: 1) We extend (Shor et al., 2024) to demonstrate self-sound-based acoustic localization using real-world acoustic data. 2) We formulate adversarial attacks on acoustic drone navigation, benchmark their effects on localization performance, and present a fully differentiable pipeline for localization and attack optimization in an acoustic setting. 3) We analyze the resource and performance implications of optimizing attack-source location. 4) We develop an acoustic-channel attack test-time defense algorithm, capable of credibly reconstructing the original perturbation waveform sampled by the drone’s acoustic sensing system.
@@ 2 Related Work
In this work we adopt the acoustic localization algorithm from (Shor et al., 2024). This algorithm performs drone localization based solely on the self-sound emitted by the drone’s propulsion system, as sampled by a circular array of microphones $\mathcal{M}$ located around the drone. The authors propose a three-step pipeline. First, a forward model is used to model the self-sound emitted by the drone’s propulsion system in free space, using a fixed, parametrized set $\mathcal{S}$ of point sound sources optimized to fit an actual free-space recording. Second, a neural, transformer-based inverse model is used to regress the drone location based on the propulsion sound as sampled by the microphone array. This sound is simulated using both the forward model and RIR by superimposing the direct path from each point sound source to each microphone, along with the reflected paths from the walls. These paths are given by the image source model (ISM), replacing source reflections on walls with imaginary point sound sources (termed images (Allen & Berkley, 1979)). Lastly, the authors present the optimization of the time-dependent angular offsets of the rotors, termed phase modulation, in order to improve localization accuracy. We further elaborate on the concept of phase modulation in Section 3.4 in the sequel. Our reason for using the approach from (Shor et al., 2024) is the fact that it presents an accurate, purely acoustic-based localization model for us to evaluate under acoustic adversarial perturbations, and mainly since the proposed phase modulation mechanism will serve us in establishing our proposed defense method in Section 3.4.
One shortcoming of (Shor et al., 2024) is the choice to compute input to the location regressor (a.k.a inverse model), i.e. the sound sampled at each sensor using a simulation superimposing a set of point sources given by the ISM, with Equation 1 being applied over each point sound source separately. While this method is relatively accurate in sufficient point source sample density (Scheibler et al., 2018), the computational costs entailed in computing the ISM render this method prohibitively expensive for larger and more complex acoustic environments. Our specific use-case of adversarial attack optimization calls for the optimization of sound emitted by the attacker (Section 3.3), obligating rapid inference of the forward model (i.e., through sound generation). This makes the original ISM-based approach of sound computation inapplicable for our purposes. In this work we therefore opt develop a modified version of the ISM-reliant algorithm from (Shor et al., 2024), making use of neural-acoustic fields (NAFs) (Luo et al., 2022), as further elaborated on Section 3.2.
@@ 3 Method
In the following section, we outline our approach to developing adversarial attacks and defenses for the task of acoustic drone localization. Section 3.1 lays the foundations of acoustic sensing-guided localization, and Section 3.2 presents the regressor that will be subjected to the adversarial attacks. Section 3.3 formulates the proposed acoustic adversarial attack, while Section 3.4 proposes a defense method relying on active rotor phase modulation.
@@ 3.1 Acoustic Localization Background
Acoustic localization is the task of identifying the location $\mathbf{x}_{\mathrm{s}}\in\mathbb{R}^{n}$ of a sound source emitting a sound signal $s_{\mathrm{s}}(t)\in\mathbb{R}^{T}$ in a known environment, based on the response of that signal $s_{\mathrm{m}}(t)\in\mathbb{R}^{T}$ as measured at a given sensor (microphone) location $\mathbf{x}_{\mathrm{m}}\in\mathbb{R}^{n}$. Here, the temporal signals are assumed to be sampled at $T$ discrete time points, and $n$ denotes the number of location degrees of freedom. While, in the most general case, $n=6$ (for location and orientation in 3D space), in this work, we focus on 2D localization with known orientation ($n=2$). State-of-the-art acoustic localization solutions (Baron et al., 2019; Shor et al., 2024) typically employ various neural architectures $\mathcal{F}(s_{\mathrm{m}}(t))$ optimized to regress $\mathbf{x}_{\mathrm{s}}$.
$s_{\mathrm{m}}(t)$ is dependent of the decay and scatter $s_{\mathrm{s}}(t)$ undergoes when propagated through the room to $\mathbf{x}_{\mathrm{m}}$. This propagation is modeled by the room impulse response (RIR) (Borish, 1984) $r(t;\mathbf{x}_{\mathrm{s}},\mathbf{x}_{\mathrm{m}},\zeta)$ – a temporal function also dependent on the source and sensor locations, as well the room geometry and physical properties, collectively denoted as $\zeta$. This function essentially describes the response in time of an impulse at $\mathbf{x}_{\mathrm{s}}$ as perceived at $\mathbf{x}_{\mathrm{m}}$ after propagating in the environment $\zeta$. The sampled sound $s_{\mathrm{m}}$ can finally attained by convolving the RIR with the input sound,
@@ 3.2 Clean localization
To circumvent the reliance on computationally complex ISMs used in (Shor et al., 2024) for RIR computation (as detailed in Section 2), we replace usage of ISMs with neural acoustic fields (NAFs) (Luo et al., 2022). NAFs offer a reliable neural representation of the RIR, and are trained over actual RIR sampled from actual room acoustics. Unlike the original forward model from (Shor et al., 2024), NAFs produce RIRs already encompassing higher-order reflection sources. Therefore, for every point sound source, modeling the drone sound $s_{\mathrm{s}}$, the entire sound waveform sampled at a given microphone, $s_{\mathrm{m}}$, can be directly derived from Equation 1 in a single forward pass, with the RIR component attained by a single query of the NAF model. This makes the approach proposed in (Shor et al., 2024) significantly faster and more reliable.
Figure 2: Adversarial Pipeline Overview - sound from both drone sound and adversarial source is convolved with NAF-induced RIRs and superimposed at the sensor, prior to being fed into the localization model. Localization loss gradients guide source amplitude optimization.
In our altered formulation for (Shor et al., 2024), for every microphone $\mu\in\mathcal{M}$, the response of each point sound-source $s_{i}\forall i\in\mathcal{S}$ as perceived at $\mu$, is computed using Equation 1 (with NAF RIRs). The total sampled sound at sensor $\mu$ is given by $s_{\mu}=\sum_{i}s_{i}$. These inputs are fed into the inverse model transformer-encoder architecture from (Shor et al., 2024) to train our ”clean” regressor, to be later evaluated under the presence of adversarial acoustic perturbations. We report clean localization performance in Section 4.2.
We emphasize that these modifications of the original localization algorithm fall outside the main focus and intended contributions of this paper, and are primarily a stepping stone towards a fast, differentiable model for the computation of $s_{\mathrm{m}}$. Nonetheless, we still deem this contribution notable.
@@ 3.3 Acoustic Adversarial Attacks
$\max_{p}\mathcal{L}(\mathcal{M}(x+p)\quad\text{s.t.}\quad\mathcal{L}_{q}(p)\leq B$ (2)
@@ 3.3.1 Perturbation Formulation
In our setting we wish to develop adversarial attacks originated from a stationary sound source located at an arbitrary location in the environment, controlled by the attacker. The motivation for this choice is that unlike common adversarial perturbations known from the field of Computer-Vision, such as 2D perturbations added to the input of image classifiers, in our setting the attacker cannot directly control the perturbation added to the model’s input signal. This is because, unlike spatial signals (e.g. 2D images), sound signal are not stationary, meaning that the presence of an acoustic perturbation at one location in the environment inevitably affects the sampled sound at another location. For this reason, we also focus our discussion in this paper on universal adversarial attacks, rather than optimizing a distinct perturbation for every drone location and orientation.
We choose to model the sound emitted by the perturbation sound source (namely, the attacker) $s_{\mathrm{p}}\in\mathbb{R}^{T}$ using a basis =$\mathcal{B}=\{f_{k}\}_{k\in K}$ of sine waves from a chosen frequency set $K$. $f_{k}\forall k\in K$ is a sine wave of frequency $k$ sampled at $T$ timesteps . The corresponding perturbation is given by spanning this basis with a set of $|K|$ learnable amplitudes $\{\alpha_{k}\}_{k\in K}$:
@@ 3.3.2 Attack Constraints
Our goal in this section is to develop a set of constraints that would allow the attack to harm the regressor’s performance while still not be easily detectable by a human or the agent. Contrary to the computer-vision domain and similar counterpart adversarial settings, where appropriate constraints are well-studied and well-established (Croce et al., 2020) (mostly $L_{\infty},L_{2}$ and $L_{1}$ norms, among others), the relatively under-explored domain of acoustic attacks does not currently hold a set of widely accepted attack constraints. We deem the acoustic setting more challenging in this aspect as, unlike vision where many scenarios may bound together under relation with human-perception, the conditions where an acoustic perturbation can be deemed feasible or non-trivial for detection vastly change under different factors (sensors, “clean” sound properties and task, among others). For this reason, we opt to propose our own constraints for the optimized adversarial perturbation, despite the fact that few other studies do address acoustic adversarial perturbation (Du et al., 2020; Bhattacharya et al., 2020). The effect of different constraint value choices over performance is studied in Section 4.2.
Frequency constraints. Our first observation in this context is that the sound emitted by the drone’s propulsion system is periodic. If the period of perturbative sound substantially differs from that of the drone, the adversarial attack could be trivially detected by the drone system. Therefore, we filter our sine basis $\mathcal{B}$ to include only frequencies integerly intertwined in the drone’s cycle from (Shor et al., 2024). The bandwidth of perturbation source-emitted frequencies is also subjected to frequency constraints given by the speaker itself. We, therefore, also clamp the bandwidth of $\mathcal{B}$ according to some minimal and maximal values (set to a minimum of $50$ Hz and a maximum of $2$ KHz in all of our experiments). All frequency-related constraints are imposed by the construction of the perturbation as detailed in Section 3.3.1.

### v2

# arXiv:2502.20325v2[cs.SD] 03 Mar 2026

@@ Abstract
Multi-rotor aerial vehicles (drones) are increasingly deployed across diverse domains, where accurate navigation is critical. The limitations of vision-based methods under poor lighting and occlusions have driven growing interest in acoustic sensing as an alternative. However, the security of acoustic-based localization has not been examined. Adversarial attacks pose a serious threat, potentially leading to mission-critical failures and safety risks. While prior research has explored adversarial attacks on vision-based systems, no work has addressed the acoustic setting. In this paper, we present the first comprehensive study of adversarial robustness in acoustic drone localization. We formulate white-box projected gradient descent (PGD) attacks from an external sound source and show their significant impact on localization accuracy. Furthermore, we propose a novel defense algorithm based on rotor phase modulation, capable of effectively recovering clean signals and mitigating adversarial degradation. Our results highlight both the vulnerability of acoustic localization and the potential for robust defense strategies.
@@ 1 Introduction
Multi-rotor autonomous aerial vehicles (drones) have seen rapid adoption across a wide range of industries (Ayamga et al., 2021; Merkert and Bushell, 2020; Moshref-Javadi and Winkenbach, 2021), including emergency medicine (Johnson et al., 2021; Zailani et al., 2020), sustainability (Dutta and Goswami, 2020; Mahroof et al., 2021), and disaster control (Daud et al., 2022; Ishiwatari, 2024), as well as recreational and commercial applications (Tan et al., 2021; Benarbia and Kyamakya, 2021). Their popularity stems from the ability to operate in challenging or hazardous environments while reducing human risk and cost.
Successful drone deployment relies on accurate and efficient navigation. This has driven major advances in localization methods (Kang et al., 2015; Dijkshoorn, 2012; Sorbelli et al., 2018), particularly with deep learning (Yousaf et al., 2022; Bisio et al., 2021; Zhang et al., 2022). Current solutions primarily depend on GPS with inertial systems, visual odometry, or active sensors such as LiDAR (Arafat et al., 2023; Aburaya et al., 2024; Niu et al., 2022; Debeunne and Vivet, 2020). While effective in many scenarios, these methods often fail under degraded lighting, occlusions, or structural constraints, and may be impractical in GPS-denied environments (Arafat et al., 2023; Al-Radaideh and Sun, 2021; Dreissig et al., 2023; Meles et al., 2023). To address these limitations, researchers have explored alternative modalities, notably acoustic localization (He et al., 2023; Sun et al., 2023; Serussi et al., 2024), which leverages sound propagation for robust localization even under visual or RF impairments (Famili et al., 2022).
@@
With the growing reliance on drones, security concerns intensify. Adversarial attacks have been shown to manipulate perception systems, leading to severe outcomes such as navigation failure or collisions (Mynuddin et al., 2023; Wisniewski et al., 2024; Guesmi and Shafique, 2024; Fu et al., 2021). While defenses against such attacks exist (Mynuddin et al., 2024; Wang et al., 2023), they have focused exclusively on visual or LiDAR-based sensing. Despite the rapid adoption of acoustic localization, its adversarial robustness remains unexplored.
To address this gap, we present the first formulation and analysis of adversarial attacks targeting acoustic drone localization. We study white-box attacks from a single omni-directional sound source external to the drone. Additionally, we leverage the phase modulation mechanism proposed in Serussi et al. (2024) to design a novel defense method that separates the clean signal from adversarial interference under minimal assumptions. We further extend Serussi et al. (2024) to real-world acoustic recordings, beyond simulated settings, enabling realistic evaluation. While our experiments focus on drone navigation, the proposed approach generalizes to any single-agent acoustic localization system.
Our main contributions are:
  1. 1.
    We formulate and benchmark adversarial attacks on acoustic localization, with a fully differentiable attack pipeline.
  2. 2.
    We analyze the computational and performance implications of perturbation source-location optimization.
  3. 3.
    We develop an acoustic-channel attack test-time defense algorithm, capable of credibly reconstructing the original perturbation waveform sampled by the drone.
  4. 4.
    We extend Serussi et al. (2024) to demonstrate self-sound-based localization using real acoustic data.
@@ 2 Related Work
In this work we adopt the acoustic localization algorithm from Serussi et al. (2024). This algorithm performs drone localization based solely on the self-sound emitted by the drone’s propulsion system, as sampled by a circular array of microphones $\mathcal{M}$ located around the drone. The authors propose a three-step pipeline. First, a forward model is used to model the self-sound emitted by the drone’s propulsion system in free space, using a fixed, parametrized set $\mathcal{S}$ of point sound sources optimized to fit an actual free-space recording. Second, a neural, transformer-based inverse model is used to regress the drone location based on the propulsion sound as sampled by the microphone array. This sound is simulated using both the forward model and RIR by superimposing the direct path from each point sound source to each microphone, along with the reflected paths from the walls. These paths are given by the image source model (ISM), replacing source reflections on walls with imaginary point sound sources (termed images - Allen and Berkley (1979)). Lastly, the authors present the optimization of the time-dependent angular offsets of the rotors, termed phase modulation, in order to improve localization accuracy. We further elaborate on the concept of phase modulation in Section 3.4 in the sequel. Our reason for using the approach from Serussi et al. (2024) is the fact that it presents an accurate, purely acoustic-based localization model for us to evaluate under acoustic adversarial perturbations, and mainly since the proposed phase modulation mechanism will serve us in establishing our proposed defense method in Section 3.4.
One shortcoming of Serussi et al. (2024) is the choice to compute input to the location regressor (a.k.a inverse model), i.e. the sound sampled at each sensor using a simulation superimposing a set of point sources given by the ISM, with equation 1 being applied over each point sound source separately. While this method is relatively accurate in sufficient point source sample density (Scheibler et al., 2018), the computational costs entailed in computing the ISM render this method prohibitively expensive for larger and more complex acoustic environments. Our specific use-case of adversarial attack optimization calls for the optimization of sound emitted by the attacker (Section 3.3), obligating rapid inference of the forward model (i.e., through sound generation). This makes the original ISM-based approach of sound computation inapplicable for our purposes. In this work we therefore opt develop a modified version of the ISM-reliant algorithm from Serussi et al. (2024), making use of neural-acoustic fields (NAFs) (Luo et al., 2022), as further elaborated on Section 3.2.
@@ 3 Method
In the following section, we outline our approach to developing adversarial attacks and defenses for the task of acoustic drone localization. Section 3.1 lays the foundations of acoustic sensing-guided localization, and Section 3.2 presents the "clean" (non-perturbed) localization setting studied in this paper. Section 3.3 formulates our threat model and the proposed acoustic adversarial attack, and Section 3.4 proposes a defense method relying on active rotor phase modulation.
@@ 3.1 Acoustic Localization Background
Acoustic localization is the task of identifying the location $\mathbf{x}_{\mathrm{s}}\in\mathbb{R}^{n}$ of a sound source emitting a sound signal $s_{\mathrm{s}}(t)\in\mathbb{R}^{T}$ in a known environment, based on the response of that signal $s_{\mathrm{m}}(t)\in\mathbb{R}^{T}$ as measured at a given sensor (microphone) location $\mathbf{x}_{\mathrm{m}}\in\mathbb{R}^{n}$. Here, the temporal signals are assumed to be sampled at $T$ discrete time points, and $n$ denotes the number of location degrees of freedom. While, in the most general case, $n=6$ (for location and orientation in 3D space), in this work, we focus on 2D localization with known orientation ($n=2$). State-of-the-art acoustic localization solutions (Baron et al., 2019; Serussi et al., 2024) typically employ various neural architectures $\mathcal{F}(s_{\mathrm{m}}(t))$ optimized to regress $\mathbf{x}_{\mathrm{s}}$.
$s_{\mathrm{m}}(t)$ is dependent on the decay and scatter $s_{\mathrm{s}}(t)$ undergoes when propagated through the room to $\mathbf{x}_{\mathrm{m}}$. This propagation is modeled by the room impulse response (RIR) (Borish, 1984) $r(t;\mathbf{x}_{\mathrm{s}},\mathbf{x}_{\mathrm{m}},\zeta)$ – a temporal function also dependent on the source and sensor locations, as well the room geometry and physical properties, collectively denoted as $\zeta$. This function essentially describes the response in time of an impulse at $\mathbf{x}_{\mathrm{s}}$ as perceived at $\mathbf{x}_{\mathrm{m}}$ after propagating in the environment $\zeta$. The sampled sound $s_{\mathrm{m}}$ can finally be attained by convolving the RIR with the input sound,
@@ 3.2 Clean localization
To circumvent the reliance on computationally complex ISMs used in Serussi et al. (2024) for RIR computation (as detailed in Section 2), we replace usage of ISMs with neural acoustic fields (NAFs, Luo et al. (2022)). NAFs offer a reliable neural representation of the RIR, and are trained over actual RIR sampled from actual room acoustics. Unlike the original forward model from Serussi et al. (2024), NAFs produce RIRs already encompassing higher-order reflection sources. Therefore, for every point sound source, modeling the drone sound $s_{\mathrm{s}}$, and the entire sound waveform sampled at a given microphone, $s_{\mathrm{m}}$, can be directly derived from equation 1 in a single forward pass, with the RIR component attained by a single query of the NAF model.
Figure 2: Adversarial Pipeline Overview - sound from both drone sound and adversarial source is convolved with NAF-induced RIRs and superimposed at the sensor, prior to being fed into the localization model. Localization loss gradients guide the optimization of source-emitted waveforms.
In our altered formulation for Serussi et al. (2024), for every microphone $\mu\in\mathcal{M}$, the response of each point sound-source $s_{i}\forall i\in\mathcal{S}$ as perceived at $\mu$, is computed using equation 1 (with NAF-produced RIRs). The total sampled sound at sensor $\mu$ is given by $s_{\mu}=\sum_{i}s_{i}$. These inputs are fed into the inverse model (transformer-encoder architecture from Serussi et al. (2024)) to train our "clean" regressor, to be later evaluated under the presence of adversarial acoustic perturbations. We report clean localization performance in Section 4.2. These modifications fall outside the main intended contributions of this paper and serve as a stepping stone toward a fast, differentiable model for the computation of $s_{\mathrm{m}}$.
@@ 3.3 Acoustic Adversarial Attacks
$\max_{p}\mathcal{L}(\mathcal{M}(x+p))\quad\text{s.t.}\quad\mathcal{L}_{q}(p)\leq B$ (2)
@@ 3.3.1 Perturbation Formulation
In our setting we wish to develop adversarial attacks originated from a stationary, omni-directional sound source located at an arbitrary location in the environment, controlled by the attacker. The motivation for this choice is that unlike common adversarial perturbations known from the field of Computer-Vision, such as 2D perturbations added to the input of image classifiers, in our setting the attacker cannot directly control the perturbation added to the model’s input signal. This is because, unlike spatial signals (e.g. 2D images), acoustic signals are not stationary – the presence of an acoustic perturbation at one location in the environment inevitably affects the sampled sound at another location. For this reason, we also focus our discussion in this paper on universal adversarial attacks, rather than optimizing a distinct perturbation for every drone location and orientation.
We choose to model the sound emitted by the perturbation sound source (namely, the attacker) $s_{\mathrm{p}}\in\mathbb{R}^{T}$ using a basis $\mathcal{B}=\{f_{k}\}_{k\in K}$ of sine waves from a chosen frequency set $K$. $f_{k}\forall k\in K$ is a sine wave of frequency $k$ sampled at $T$ timesteps . The corresponding perturbation is given by spanning this basis with a set of $|K|$ learnable amplitudes $\{\alpha_{k}\}_{k\in K}$:
@@ 3.3.2 Attack Constraints
Our goal in this section is to develop a set of constraints that would allow the attack to efficiently harm localization performance while still not be easily detectable by the localizing agent. Contrary to the computer-vision domain and similar counterpart adversarial settings, where appropriate constraints are well-studied and well-established (see Croce et al. (2020)) (mostly $L_{\infty},L_{2}$ and $L_{1}$ norms, among others), the relatively under-explored domain of acoustic attacks does not currently hold a set of widely accepted attack constraints. We deem the acoustic setting more challenging in this aspect as, unlike vision where many scenarios may bound together under relation with human-perception, the conditions where an acoustic perturbation can be deemed feasible or non-trivial for detection vastly change under different factors (e.g. sensors, "clean" sound properties and task). For this reason, we propose our own constraints for the optimized adversarial perturbation. The effect of different constraint value choices over performance is studied in Section 4.2.
Frequency constraints. Our first observation in this context is that the sound emitted by the drone’s propulsion system is periodic. If the period of perturbative sound substantially differs from that of the drone, the adversarial attack could be trivially detected by the agent. Therefore, we filter our sine basis $\mathcal{B}$ to include only frequencies integerly intertwined in the drone’s cycle. The bandwidth of perturbation source-emitted frequencies is also subjected to frequency constraints given by the sound-source itself. We, therefore, also clamp the bandwidth of $\mathcal{B}$ according to some plausible minimal and maximal values (set to a minimum of $50$ Hz and a maximum of $2$ KHz in all of our experiments). All frequency-related constraints are imposed by the construction of the perturbation as detailed in Section 3.3.1.